OpenStudy (anonymous):

The value of y varies directly with x. if y=80 when x is 16, what is y when x is 80? I will give a medal. Please help solve with me.

OpenStudy (jdoe0001):

\(\bf \begin{array}{cccllll} \textit{something }&\textit{varies directly to }&\textit{something else}\\ \quad \\ \textit{something }&={\color{red}{ \textit{some value }}}&\textit{something else}\\ \quad \\ y&={\color{red}{ n}}&x&\implies y={\color{red}{ n}}x \end{array}\\ \quad \\ y=80\qquad x=16\qquad y={\color{red}{ n}}x\implies 80={\color{red}{ n}}\cdot 16\) solve for "n", or the "constant of variation" then plug it back in the equation to know what "y" is when x = 80 just set in the equation x to 80 :)

OpenStudy (anonymous):

5*16=80

OpenStudy (jdoe0001):

well... you mean n = 5 yes so that means the equation is y = 5x what's "y" if x = 80? well \(\bf y=5x\qquad x=80\implies y=5\cdot 80\)

OpenStudy (anonymous):

y=400

OpenStudy (jdoe0001):

yeap since n=5 \(\bf y=80\qquad x=16\qquad y={\color{red}{ n}}x\implies 80={\color{red}{ n}}\cdot 16\implies \cfrac{\cancel{ 80}}{\cancel{ 16}}=n=5\) thus y = 5x

OpenStudy (anonymous):

Thanks, thats the final answer right?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

yw